# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

### v2

@@ 2. Semantic Operators
3. The LOTUS Programming Interface
@@
In this section we describe the LOTUS API, which implements the semantic operator model as an extension of Pandas (pan, [n. d.]). We chose a Pandas-like API in our initial system implementation to make it easy for users to integrate LOTUS with popular AI libraries in Python. However, semantic operators could also be added to a variety of other data processing APIs and query languages, such as SQL.
@@ 3.1. Datatypes
3.1.1. Data Model
LOTUS’ data model consists of tables with structured and unstructured text fields. LOTUS’ semantic operators can take both of these data-types as inputs. For instance, Figure 2 shows an example LOTUS program, where the table schema over ArXiv papers consists of attributes for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date. These columns are then passed as parameters to semantic operators, such as sem_search, sem_filter, sem_agg.
Additionally, LOTUS supports semantic indices over natural-language text columns to provide optimized query processing. These indices leverage semantic embeddings over each document in the column to capture semantic similarity using embedding distance metrics. Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded using load_sem_index.
3.1.2. Parameterized Natural Language Expressions (langex)
Programmers write each langex in natural language, parameterized by one or more data columns, which are indicated in the brackets within a string, or double brackets within a formatted string, as Figure 2 shows. Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user. Operator-specific LLM prompts based on these expressions are automatically constructed by LOTUS’ underlying query engine, and the underlying system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
@@ 3.2. Semantic Operators in LOTUS
We now overview the semantic operators supported in the LOTUS API, which includes the core set of operators described in Section 2, as well as several additional variants provided for convenience.
Sem_filter, Sem_join, & Sem_sim_join. The LOTUS API supports sem_filter and sem_join, both of which take a langex predicate, as described in Section 2. In addition, LOTUS provides a join variant, sem_sim_join, where tuples are matched according to their semantic similarity, rather than an arbitrary natural-language predicate. Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine. Figure 3 provides an example of the sem_join compared to the sem_sim_join, where the user specifies the left and right table join keys, and a parameter .
@@
Sem_agg performs an aggregation over the input relation, with as langex signature that provides a commutative and associative aggregation function, as shown by Figure 2. Similar to sem_topk, LOTUS allows users to specify a group-by parameter.
@@
Sem_group_by creates groups over the input dataframe according to the langex grouping criteria by clustering the data into a target number of groups (e.g., papers_df.sem_group_by("the topic of the abstract", 20)). By default, the operator automatically selects group labels, but the user can also optionally specify target labels. This operator is useful both for unsupervised discovery of semantic groups, and for semantic classification tasks.
@@
Sem_map & Sem_extract both perform a natural language projection over an existing column. While sem_map projects to an arbitrary text attribute, sem_extract projects each tuple to a list of sub-strings from the source text. This is useful for applications, such as entity extraction or fact-checking, where finding snippets or verified quotes may be preferable to synthesized answers.
4. Execution Plans and Optimizations for Semantic Operators
Semantic operators create a rich design space of diverse execution plans which have significant consequences on system efficiency and accuracy. For example, there are many possible algorithms to join two collections of records using a LLM and a natural language expression (all the standard join algorithms, as well as ML heuristics to approximate the join, e.g., by semantically clustering similar records using an embedding model).
Specifically our goal is to offer an efficient plan while while guaranteeing high accuracy. Defining accuracy for semantic transformations, however, presents a fundamental challenge in the absence of ground truth labels. Therefore, we define accuracy of each operator plan with respect to a ”gold” algorithm, which uses a powerful oracle model . These ”gold” algorithms are tractable solutions that can achieve high quality results, but may be expensive in terms of execution time and LM call complexity. We discuss the design space of these ”gold” algorithms for several core semantic operators we introduced. In addition, we present several novel optimizations for semantic filters, joins, top-k and group-by. These optimizations are either lossless or provide statistical guarantees on accuracy with respect to a ”gold” algorithm with the expensive model . Specifically, our approximations emulate the results of the expensive but high-quality algorithm by using model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) or semantic indexes, building on past work on accelerating ML-based query processing with guarantees (Kang et al., 2022, 2019) but achieving this property for new types of operators. Notably, these optimizations can all be performed declaratively given a user’s accuracy targets. We show in Section 5 that they significantly accelerate real applications.
While this paper proposes new optimizations for only a few, expensive semantic operators, we envision that a wide range of other new and traditional optimizations are possible under the semantic operator model. For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g. operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g. code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
4.1. sem_filter
4.1.1. Gold Algorithm
Our gold algorithm for the semantic filter runs batched LLM calls over the tuples in relation . Each tuple, , is processed independently using the model , which is prompted to output a boolean value indicating whether the tuple passes the langex predicate.
4.1.2. Algorithmic Approximation
We can declaratively optimize semantic filters by providing an approximation which will obtain recall and precision targets and with respect to the gold implementation with probability . To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which provides a score indicating whether the tuple passes the predicate, in conjunctions with the oracle LLM . This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings. Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms. Additionally, other works, which study cascades for LLMs, use heavy-weight scoring mechanisms to decide whether to use the proxy model, which incurs a heavy execution time penalty, and does not provide accuracy guarantees. In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
Providing accuracy guarantees is important because, in general, we cannot assume the proxy model is accurate. In fact, the proxy model, may perform very poorly for some tasks and datasets. The goal of our algorithm is to automatically discover the quality of the proxy by sampling and comparing to the gold algorithm with the oracle model. If the proxy provides high quality results, our algorithm will exploit it, leading to lower execution time. However, if the proxy model performs poorly, we will selectively route to it according to the proxy scores , and a learned threshold on the proxy scores. As prior work (Kang et al., 2020) discusses, such sampling-based algorithms to learn a threshold, , introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
Our algorithm leverages a small LLM as the proxy model. We generate scores using the the log-probabilities corresponding to the True or False output tokens of the proxy model’s answer and re-scaling by the quantiles over the proxy scores of tuples in the relation.
Algorithm 1 shows the full procedure for performing the approximate semantic filter. We begin by using an collecting a sample , which we label with the oracle and proxy models. The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022). Using a the central limit theorem and the normal approximation on the distribution of sample statistics, we then find thresholds and , which will respectively ensure the desired precision target and recall target are met, each with an error probability of . These sub-procedures follow the statistical techniques used in prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
Finally the algorithm proceeds to process each tuple in the relation. If the tuple’s proxy score is at least , we mark it as passing the predicate. If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
4.2. sem_join
4.2.1. Gold Algorithm
Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left table, and right table, . The gold algorithm implements a nested-loop join pattern with efficient batched inference to maximize GPU utilization. This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively. Each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate. This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
4.2.2. Algorithmic Approximation
Similar to semantic filters, we declaratively optimize semantic joins by providing an approximation that obtains recall and precision targets and with respect to the gold algorithm with probability . Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we leverage an even cheaper proxy based on semantic similarity scores, which can be obtained efficiently with a semantic index over the join keys. We leverage two possible join approximation algorithms and dynamically chooses the lowest cost one, which will minimize the number of LLM calls.
The first approximation, sim-filter produces a proxy score , based on embedding similarity. We leverage the semantic index to perform batched similarity search and re-calibrate similarity scores according to the quantiles of the distribution over all pairs to obtain proxy scores. This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate. This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024) and is not always present. Thus, we introduce an alternative approximation suitable when correlation is not present.
The second approximation, map-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and . As before we also re-calibrate the proxy scores taking their quantiles. Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity. The projection step invokes the LM over each tuple in left join table, prompting it to predict values in the attribute domain of the right join key that match. Notably, this LM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation, which can optionally take user-provided demonstrations. Figure 5 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset. Here, the map step would invoke the LM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters. We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample. We then obtain , and independently for each approximation algorithm using their respective proxy scores. Using the tuned thresholds and known proxy scores over for both algorithms over the dataset, we then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
4.3. sem_group_by
Performing this semantic operator entails discovering group labels and classifying each input tuple using the discovered labels. In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n. d.]). Our gold algorithm instead offers a tractable implementation to obtain group labels using a linear pass over the data with the reference model . Furthermore, once the labels are discovered, we perform the LM-based classification task and show how to optimize this step using an approximation with probabilistic guarantees.
4.3.1. Gold Algorithm
Our semantic group-by algorithm proceeds in two stages. The first stage performs the unsupervised label discovery, given the user’s grouping criteria and target number of groups. This stage first performs a semantic projection, prompting the LM to predict a label for each input tuple. Then, we create a semantic index over the projected column and perform a vector clustering over these predicted labels, using FAISS’ efficient k-means implementation to construct groups. Lastly, for each group, we top-k sample documents based on their similarity to the centroid, and few-shot prompt the LM to generate a label describing the given documents according to the user-provided grouping criteria (e.g. the topics of the documents) using a semantic aggregation.
In the second stage of the algorithm, we use the generated labels to perform classification over each document. We do this using a semantic map, which prompts the LM to output a label conditioned on the candidate labels and document.
4.3.2. Algorithmic Approximation
We provide an efficient approximation, specifically for the classification task in the second stage of the algorithm. Similar to semantic filters and join, we perform this optimization declaratively, while guaranteeing a classification accuracy target is met with probability . Similar, to semantic joins, we leverage semantic similarity scores as a cheap proxy. Specifically, we leverage the index constructed in stage one of the algorithm and compute similarity scores between the embeddings of each document’s un-grounded prediction and the embeddings of each class label. The proxy score of the tuple with a predicted label, , for group label is given by . Our goal is then to learn a threshold, , such that if the similarity between an embedding and the class label’s embedding is greater than , we return the class label, and otherwise we resort to the more expensive LM-based classification procedure. We accomplish this by uniform sampling and running the sub procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
4.4. sem_topk
Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data. While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n. d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, our implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low execution time. Here we outline leverage several algorithmic design decisions for the gold algorithm as well as a lossless optimization.
4.4.1. Gold Algorithm
Two important algorithmic design decisions for the semantic top-k include how to implement the LM-based comparison and how to aggregate ranking information across LM-based comparisons. Our implementation uses pairwise LM-based comparisons for the former, and a quick-select based top-k algorithm (Hoare, 1961) for the latter.
We briefly describe the reason for these choices and alternatives considered. Pairwise-prompting methods offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label. The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020). In contrast, pairwise comparisons have been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
In addition, we studied several alternative rank-aggregation algorithm, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm, which we discuss in Section 5. The quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot. Because each round is fully parallelizable, we can efficiently batch these LM-based comparisons before recursing in the next round. In section 5, we find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation.
4.4.2. Optimizations
We leverage the semantic index to optimize pivot selection for some queries, while incurring no accuracy loss. This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores. In this case, LOTUS can sort the document set based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot. This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss. In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
4.5. sem_agg
Performing semantic aggregations is inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows. Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length. LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results.
4.5.1. Gold Algorithm
Our implementation builds on the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n. d.]; lan, [n. d.]). These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains. Our aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024).

### v3

@@ 2.3. Core Semantic Operators
Gold Algorithm. Since this operator, by definition, entails a clustering problem, known to be NP-hard (Dasgupta, [n.d.]), our gold algorithm instead offers a computable implementation to obtain group labels using a tractable clustering algorithm to discover group labels followed by a point-wise classification step. Both steps use a linear LLM pass over the relation.
Specifically, the first stage performs an LLM-based clustering algorithm to discover centers , given user langex. Our gold algorithm first performs a semantic projection, prompting the LLM to predict a label for each input tuple. Then, we embed these candidate labels and perform an efficient vector clustering using a k-means algorithm to construct groups. For each group, we top-k sample tuples based on centroid-similarity scores, and use a semantic aggregation that prompts the model to generate an appropriate label over each group. In the second stage of our gold algorithm, we use the generated labels, and invoke the model to perform point-wise assignments .
3. Optimized Execution Plans for Semantic Operators
In this section, we present novel optimizations for several costly operators, including semantic filters, joins, top-k and group-by. Each optimization provides statistical accuracy guarantees with respect to the gold algorithm, building from Section 2. While this paper focuses on novel optimizations for several expensive semantic operators, we envision a rich space of future work exploring new optimizations and applications of traditional optimizations to semantic operator programs. For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g. operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g. code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
3.1. Optimizing Semantic Filter
We provide an approximation for semantic filters that obtains recall and precision targets and with respect to the gold implementation with probability . To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which outputs a score indicating whether the tuple passes the predicate. This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings. Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms. Other works study cascades for LLMs, but use heavy-weight scoring mechanisms to decide whether to use the proxy model and do not provide accuracy guarantees. In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
In general, we cannot assume the proxy model is accurate. In fact, the proxy model may perform very poorly for some tasks and records. The goal of our algorithm is to automatically discover the quality of the proxy at query time by sampling and comparing to the gold algorithm with the oracle model. Our algorithm will exploit the proxy when it is likely to provide high-quality outputs, and otherwise defer the oracle model. To do this we must learn a decision rule, with learned thresholds on the proxy scores using sampling. As prior work (Kang et al., 2020) discusses, such sampling-based algorithms introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
We consider a small LLM as the proxy model and generate scores using log-probabilities555This optimizations assumes access to log-probabilities, which is available in common model providers, such as OpenAI, and serving systems, such as vLLM. In the absence of available log-probabilities, alternative uncertainty quantification methods, such as prompt-based methods, may be more suitable. corresponding to the True or False output tokens of the proxy model’s predicate evaluations, re-scaling by the quantiles over all generated log-probabilities over the relation.
Algorithm 1 shows the full procedure for performing the approximate semantic filter. We begin by collecting a sample of tuples , and labeling each sample with the oracle and proxy models. The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022). Using the central limit theorem and the normal approximation on the distribution of sample statistics, we then learn a decision rule by finding thresholds and , which will respectively ensure the precision target and recall target are met, each with an error probability of . The statistical sub-procedures to choose each threshold follows prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
Finally, using the learned decision rule, our algorithm proceeds to process each tuple in the relation. If the tuple’s proxy score is at least , we mark it as passing the predicate. If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
3.2. Optimizing Semantic Join
Similar to semantic filters, we provide an approximation for semantic joins that obtains recall and precision targets and with respect to the gold algorithm with probability . Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we consider an even cheaper proxy based on semantic similarity scores using embeddings. We leverage two possible proxy algorithms and dynamically choose the lowest cost one.
The first approximation, sim-filter produces a proxy score , based on embedding similarity. We perform batched similarity search between the right and left join keys and re-calibrate similarity scores according to their quantiles to obtain proxy scores. This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate. This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024b) and is not always present. Thus, we introduce an alternative approximation, often suitable when correlation is not present.
The second approximation, project-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and . As before, we also re-calibrate the proxy scores taking their quantiles. Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity. The projection step invokes the LLM over each tuple in left join table, prompting it to predict values for the right join key. Notably, this LLM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation. Figure 3 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset. Here, the map step would invoke the LLM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters. We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample. We then obtain tresholds and independently for each approximation algorithm using their respective proxy scores, and . We use the learned thresholds for each candidate plan to then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
3.3. Optimizing Semantic Group-by
We provide an efficient approximation that guarantees a classification accuracy target, , is met with probability . We follow the clustering algorithm in the first stage of the gold algorithm to discover centers . We then use a proxy-based approximation for point wise assignments in the second stage of the algorithm. Similar, to semantic joins, we consider semantic similarity scores as a cheap proxy, although other proxy models are also feasible. Specifically, we leverage the embeddings constructed during the clustering stage and compute similarity scores between the the candidate label of each tuple and the discovered centers . The proxy score of the tuple with a predicted label, , for a discovered center is given by . Our goal is then to learn a threshold, , such that if the proxy score between a tuple and center is greater than , we return the center as label for the tuple, and otherwise we resort to the more expensive LLM-based classification procedure. We accomplish this by uniform sampling and running the sub-procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
3.4. Optimizing Semantic Top-k
We leverage the embedding similarity scores to optimize pivot selection for some queries, while incurring no accuracy loss. This optimization is useful when there exists correlation between the ranking imposed by the user’s arbitrary sorting criteria and the ranking imposed by semantic similarity scores. In this case, we can sort tuples based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot. This can reduce the number of LLM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss. In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
4. The LOTUS Systems
@@
In this section we describe the LOTUS system, which implements the semantic operator model as an extension of Pandas (pan, [n.d.]). We chose a Pandas-like API in our initial system implementation to make it easy for users to integrate LOTUS with popular AI libraries in Python. However, semantic operators could also be added to a variety of other data processing APIs and query languages, such as SQL. LOTUS leverages vLLM (Kwon et al., 2023) to perform efficient batched inference, and uses FAISS (Douze et al., 2024; Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations, with indices stored and maintained locally on disk, by default.
@@ 4.1. Datatypes
LOTUS’ data model consists of tables with structured and unstructured fields (e.g., text or images). LOTUS’ semantic operators can take both of these data-types as parameters to a langex.
Additionally, LOTUS supports semantic indices over natural-language text columns to provide optimized query processing. These indices leverage semantic embeddings over each document in the column and capture semantic similarity using embedding distance metrics. Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded from disk using load_sem_index.
@@ 4.2. Semantic Operators in LOTUS
We now overview the semantic operators supported in the LOTUS API, which includes the core set of operators described in Section 2, as well as several additional variants provided for convenience. Each operator takes optional parameters to specify a target accuracy and error probability, which the optimizer will use to transparently perform optimizations.
Sem_filter, Sem_join, & Sem_sim_join. The LOTUS API supports sem_filter and sem_join, both of which take a langex predicate, as described in Section 2. In addition, LOTUS provides a join variant, sem_sim_join, where tuples are matched according to their semantic similarity, rather than an arbitrary natural-language predicate. Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine by leveraging vector similarity search. Figure 4 provides an example of the sem_join compared to the sem_sim_join, where the user specifies the left and right table join keys, and a parameter .
@@
Sem_agg performs an aggregation over the input relation, with a langex signature that provides a commutative and associative aggregation function, as shown by Figure 2. Similar to sem_topk, LOTUS allows users to specify a group-by parameter.
@@
Sem_group_by creates groups over the input dataframe according to the langex grouping criteria and target number of groups. By default, the operator discovers suitable group labels, but the user can also optionally specify target labels. This operator is useful both for unsupervised discovery of semantic groups, and for semantic classification tasks.
@@
Sem_map & Sem_extract both perform a natural language projection over an existing column. While sem_map projects to an arbitrary text attribute, sem_extract projects each tuple to a list of sub-strings from the source text. This is useful for applications, such as entity extraction, where finding snippets or verified quotes may be preferable to synthesized answers.